505:

How do you know that 6:14 is equivalent to 15:35 using a multiplication table?

Vocaloid:

starting with 6:14 using a multiplication table, 6 = 2*3 14 = 2*7 so the ratio is now: 2*3 : 2*7 crossing out 2 from both sides 3:7 using a multiplication table, you can see that multiplying both of these numbers by 5 gives us 15 and 35 so 6:14 = 3:7 = 15:35